# Spin and energy diffusion vs. subdiffusion in disordered spin chains

### v1

@@ Abstract
While the high-temperature spin diffusion in spin chains with random local fields has been the subject of numerous studies concerning the phenomenon of many-body localization (MBL), the energy diffusion in the same models has been much less explored. We show that energy diffusion is faster at weak random fields but becomes essentially equal at strong fields; hence, both diffusions determine the slowest relaxation time scale (Thouless time) in the system. Numerically reachable finite-size systems reveal the anomalously large distribution of diffusion constants with respect to actual field configurations. Despite the exponential-like dependence of diffusion on field strength, results for the sensitivity to twisted boundary conditions are incompatible with the Thouless criterion for localization and the presumable transition to MBL, at least for numerically reachable sizes. In contrast, we find indications for the scenario of subdiffusive transport, in particular in the dynamical diffusivity response.
@@ I Introduction
So far, numerical studies of high-$T$ spin d.c. conductivity $\sigma^{0}_{s}$ and the corresponding spin diffusion constant ${\cal D}^{0}_{s}$ established exponential-like dependence on the disorder strength $W$ [16, 30, 19, 20, 31, 32], which appears consistent with the related exponential increase of the Thouless time $\tau_{\mathrm{Th}}(W)$ [10, 26]. However, due to the modest system length $L$ reachable numerically, results are heavily influenced by large random-field sample-to-sample variations [32, 28, 33]. This contrasts analogous transport results for Heisenberg chains with quasiperiodic fields [34], where fluctuations are minimal. In such systems, one can also identify MBL-like behavior (e.g., exponential-like dependence of transport properties on disordered strength $W$) [35, 36, 37, 38, 39, 40, 41, 42], being also realized experimentally in cold-atom systems [43, 44]. Recently, the debate of the MBL regime has been revived by the analytical study indicating that even at large disorder, the transport is not localized but rather subdiffusive [45], having the origin in insulating/blocking regions and related Griffiths effects already considered in MBL models [17, 46, 47].
@@ II Spin and energy diffusion
with $S=1/2$ spin operators and random fields $w_{l}\in[-W,W]$ with the uniform probability distribution. We consider the chain with $L$ sites and periodic boundary conditions numerically. To allow the Thouless approach to test transport properties via sensitivity to (twisted) boundary conditions, as well the validity of RMT universality, [53, 34] we introduce in general a finite flux $\varphi\neq 0$. As with most MBL studies, we fix for convenience in numerical studies the (Ising) interaction corresponding to the isotropic case $\Delta=1$ and use $J=1$ as the unit of energy. Within the MBL problem, we are interested in the dynamical transport properties at finite temperature $T>0$, and as usual, for simplicity, we deal with the $T\to\infty$ limit. While d.c. transport quantities (i.e., the spin conductivity $\sigma^{0}_{s}$ and thermal conductivity $\kappa^{0}$) vanish at $T\to\infty$, the related diffusion constants do not and remain highly nontrivial quantities to determine.
@@
In the last equation, ${\cal D}_{s}(\omega)$ is expressed directly in terms of spin-current matrix elements among many-body (MB) eigenstates $|n\rangle$, $|m\rangle$ and $N_{st}$ representing the total number of MB states. The spin current $j_{s}$ and high-$T$ static susceptibility $\chi_{s}^{0}=\chi^{0}_{q\to 0}=\tilde{\chi}_{s}^{0}/T$ are within the considered model (1) given by
@@
where ${\cal D}_{s}^{0}={\cal D}_{s}(\omega\to 0)$ is d.c. spin diffusion constant and $\Delta e=\overline{e_{n+1}-e_{n}}$ is the average level spacing, or equivalently, $\rho=1/\Delta e$ the MB density of states. We note that the Thouless criterion, applied to the noninteracting Anderson model [53], requires $R<1$ for localization, but as well that that $R$ decreases/vanishes with increasing $L$. It is also evident that in the finite-$L$ system with $R<1$, one generally cannot expect the validity of RMT universality since the MB levels do not vary enough with $\varphi$ to satisfy the concept of avoided crossing. Consequently, the level sensitivity $R\sim 1$ is analogous to the criterion of merging the Heisenberg vs. Thouless time $\tau_{\mathrm{H}}\sim\tau_{\mathrm{T}h}$ [10], where $\tau_{\mathrm{H}}\sim 1/\Delta e$ and $\tau_{\mathrm{T}h}\sim L^{2}/{\cal D}^{0}_{s}$.
Figure 2: Level sensitivity $R$ (averaged over random field configurations): (a) vs. $W$ for different $L$: for $L=14-18$ obtained via direct ED, and for $L=20-28$ via MCLM with approximated $\Delta e$ (see text for details), (b) the same averaged $R$ vs. $L$ for different $W$.
In Fig. 1, we present results for RMT markers for the spin current, $Y_{s}$, $Q_{s}$ and the energy current $Y_{e}$, $Q_{e}$, as well as the level sensitivity $R$, evaluated according to Eq. (9). Results are obtained via ED in the chain with $L=18$ sites for a few ($N_{c}=10$) fixed random configurations $\eta_{l}=[-1,1]$ with fields $w_{l}=W\eta_{l}$ steadily increasing with $W$. The sums over eigenstates in Eq. (3) are here performed over one-quarter of states, i.e., $\tilde{N}_{st}=N_{st}/4$ states in the middle of the MB spectrum. The presented results offer some nontrivial conclusions: (a) The variation of RMT markers vs. $W$ in Figs. 1(a,b) are quite similar for spin and energy current matrix elements. (b) The deviations for GOE values $Y_{s,e}\sim 1$ and $Q_{s,e}\sim 3$ appear in each sample quite simultaneously with the crossover at $R\sim 1$. (c) While the sample-sample fluctuations of d.c. values ${\cal D}^{0}_{s,e}$ are already substantial at intermediate $W\sim 1.5-2$ (see the discussion below), each sample still reveals $Q_{s,e}\sim 3$. This allows the conclusion that for fixed $w_{l}$ configurations, the variations of calculated d.c. values within rather broad MB spectrum $e_{n}$ (one quarter of) are quite modest.
@@ IV Distribution of diffusion constants
It has been previously numerically established, at least for the spin diffusion ${\cal D}_{s}^{0}$ [32], that results reveal a broad distribution of values in reachable systems $L\lesssim 28$. The latter cannot be captured within a simple Gaussian distribution even at intermediate $W\sim 2$, becoming closer to the log-normal distribution up to the (here numerically) reachable $W^{*}\lesssim 4$. We note that such anomalous distributions are very distinct from the Anderson model of noninteracting particles in higher dimension $D>1$, where due to large reachable sizes $L^{D}\sim 10^{7}$ [31] evidently yield already self-averaged results.
Figure 3: (a) Different presentations of the distribution of d.c. spin diffusion values ${\cal D}^{0}_{s}$: (1) vs. unrenormalized $W$, (2) using renormalized variance $\sum_{l}\eta^{2}_{l}/L=1/3$, and (3) presenting results vs. effective $\tilde{W}=W/(\bar{\eta}^{2}/3)$ (see text for details). (b) and (c) shows the distribution of results for ${\cal D}_{s}^{0}$ and ${\cal D}_{e}^{0}$, respectively, vs. effective $\tilde{W}$ obtained via MCLM for $L=24$, while the inset displays also the comparison with $L=20$ and $L=28$ results. Solid lines refer to the level sensitivity $R=1$, i.e., the estimated limit of validity of GOE. Dashed lines in panels (b,c) depict $\propto\exp(-bW/J)$ with $b=2.5$.
@@
Figure 4: Cumulative distribution function (CDF) for (a) d.c. spin diffusion ${\cal D}_{s}^{0}$, (b) d.c. energy diffusion ${\cal D}_{e}^{0}$, obtained via MCLM on $L=24$ sites for different $W=0.5-4$, and (c) the direct comparison between both for $W=1.0,2.0,3.0,4.0$, including also $L=20\,,28$ results.
@@ V Evidence for subdiffusion
Figure 5: System size $L$ dependence of spin diffusion ${\cal D}_{s}^{0}$ for $W=1,2,3$. Shown are MCLM (ED) results for $L\geq 20$ ($L\leq 18$) and three $N_{c}=3$ different configurations. The dotted lines are guides to the eye, denoting general trend with $L$.
@@
The above simple scenario for diffusion in MBL systems has implications. Namely, it should apply also to spin chains with quasiperiodic (QP) fields studied both theoretically/numerically [35, 36, 37, 38, 39, 40, 41, 42, 34], as well as experimentally [43, 44]. Indeed, similar exponential dependence on field strength $W$ has been found for ${\cal D}_{s}^{0}$ in QP systems [34]. With the essential difference that in the latter deterministic case (the absence of randomness), there are no significant (up to boundary effects) variations of ${\cal D}_{s}^{0}$ with actual field configuration. In contrast, for random but uncorrelated fields, as discussed in this study, there are large variations in $J_{l}^{\mathrm{eff}}$ resulting in variations of the effective distance between resonant islands $d$, as well as $w_{l}$ in the localized regions. These ingredients are the argument for the Griffiths scenario leading to domination and blocking of longer insulating islands [17, 46, 47, 32]. Furthermore, such arguments are also the basis for the analytical study supporting the subdiffusive transport in the thermodynamic limit [45].
Figure 6: Dynamical (a) spin and (b) energy diffusivity ${\cal D}_{s,e}(\omega)$ in the log-log scale, calculated for a single field configuration and $W=0.5-3.5$ on chains with different lengths, i.e., $L=18$ using ED, and $L=20-28$ employing MCLM.
@@
In contrast, one expects more pronounced effects in the dynamical diffusivity ${\cal D}_{s,e}(\omega)$ since results cover a much larger dynamical range, taking into account also $\omega>\delta\omega\sim 10^{-4}$. As the signature of subdiffusion one can identify the dependence ${\cal D}(\omega)\sim{\cal D}^{0}+c|\omega|^{\alpha}$ with $\alpha<1$ [17, 46] in the low $\omega<\omega^{*}$ range. Such behavior implies that even with vanishing (or very small) ${\cal D}^{0}$, the d.c. polarizability $\chi_{p}\propto\int\mathrm{d}\,\omega{\cal D}_{s}(\omega)/\omega^{2}$ [20] is diverging, in contrast to the localization defined by $\chi_{p}<\infty$ requiring $\alpha\geq 1$. Numerical investigation of finite systems indicated that ${\cal D}_{s,e}(\omega)$ [and related conductivity $\sigma_{s}(\omega)$] at $\omega\sim{\cal O}(0.1)$ is indeed consistent with $\alpha\simeq 1$ [60, 16, 65, 19]. A similar conclusion (in a similar frequency range) can be reached by analysis of the short wave-length $q\to\pi$ response (i.e., the so-called imbalance) [25, 66, 32]. Here, contrary, we show that a closer analysis of $\omega\ll 1$ (obtained with upgraded MCLM method with $\delta\omega\sim 10^{-4}$ resolution) indicates that random-field results are generally at larger $W$ consistent with $\alpha<1$. The latter contrasts the QP field results, for which $\alpha\simeq 1$ is reported down to $\omega\to 0$.
@@
Figure 7: Dynamical spin diffusivity ${\cal D}_{s}(\omega)$ (in log-log scale) for several random field configurations with few $W$, calculated on $L=28$ chain. Each panel represents a different $w_{l}$ configuration. The dashed line corresponds to the subdiffusive $\alpha=0.5$ power law.
In Fig. 6, we present numerical results for dynamical ${\cal D}_{s,e}(\omega)$, obtained for single field configurations but different $W=0.5-3.5$ on different systems sizes $L=18-28$. While $L=18$ can reached by ED directly evaluating Eq. (3), $L\geq 20$ results are obtained via MCLM with finite frequency resolution $\delta\omega\sim 10^{-4}$. Apart from fluctuations in MCLM results, being the consequence of finite $\delta\omega$ as well of restricted sampling over initial microcanonical energies ${\cal E}$, results are consistent, provided that we input the same field configurations $w_{l}$. The dependence of MCLM results on specific $w_{l}$ configurations is presented in Fig. 7 for two different field configurations. For reference, we also plot lines with characteristic $\alpha=0.5$ subdiffusion exponent. It is evident that for larger $W\geq 2.0$, we find evidence for $\alpha<1$ but with a shrinking range $\omega<\omega^{*}(W)$ with increasing $W$. As apparent from Fig. 7, for larger $W\geq 3$ the subdiffusive range $\omega^{*}(W)$ is approaching numerically limiting $\delta\omega$, while the overall $\omega>\omega^{*}$ dependence is then closer to the power law with $\alpha\lesssim 1$.
Finally, to show the difference between the QP and random-field model, we present a detailed analysis of the ${\cal D}_{s}(\omega)$ spectra. Firstly, in Fig. 8(a,b), we show (in log-log and normal scale, respectively) fits to ${\cal D}(\omega)\sim{\cal D}^{0}+c|\omega|^{\alpha}$ with different $\alpha<1$ within $\omega<\omega^{*}$ range with shrinking $\omega^{*}(W)$. On the other hand, the results for QP at low-$\omega$ are different. In Fig. 8(c) we show MCLM results for $L=27$ system with QP potential [34] $w_{l}=W\cos(ql+\phi_{0})$ with (nearly) irrational $q/(2\pi)=11/27\sim(3-\sqrt{5})/2$ and $\phi_{0}=0$. For all considered values $W$, we observe consistently linear $\omega$ dependence, i.e., $\alpha\sim 1$ also in the most relevant $\omega\ll 1$ regime.
@@
Figure 8: Dynamical spin diffusivity ${\cal D}_{s}(\omega)$, calculated via MCLM for (a,b) random-field model (in log-log and normal scale, respectively) for one field configuration and $L=24$. Dashed lines represent fits to the ${\cal D}^{0}+c|\omega|^{\alpha}$ with (top to bottom) $\alpha=0.61\,,0.58\,,0.64\,,0.83$. (c) ${\cal D}_{s}(\omega)$ for quasiperiodic system ($L=27$). See text for details.

### v2

@@ Abstract
While the high-temperature spin diffusion in spin chains with random local fields has been the subject of numerous studies concerning the phenomenon of many-body localization (MBL), the energy diffusion in the same models has been much less explored. We demonstrate that energy diffusion is faster at weak random fields but becomes essentially equal at strong fields; hence, both diffusions determine the slowest relaxation time scale (Thouless time) in the system. Numerically reachable finite-size systems reveal the anomalously large distribution of diffusion constants with respect to actual field configurations. Despite the exponential-like dependence of diffusion on field strength, the results for sensitivity to twisted boundary conditions are incompatible with the Thouless criterion for localization and the presumed transition to MBL, at least for numerically reachable sizes. In contrast, we find indications of the scenario of subdiffusive transport, particularly in the dynamical diffusivity response.
@@ I Introduction
So far, numerical studies of high-$T$ spin d.c. conductivity $\sigma^{0}_{s}$ and the corresponding spin diffusion constant ${\cal D}^{0}_{s}$ established exponential-like dependence on the disorder strength $W$ [16, 30, 19, 20, 31, 32], which appears consistent with the related exponential increase of the Thouless time $\tau_{\mathrm{Th}}(W)$ [10, 26]. However, due to the modest system length $L$ reachable numerically, results are heavily influenced by large random-field sample-to-sample variations [32, 28, 33]. This contrasts with analogous transport results for Heisenberg chains with quasiperiodic fields [34], where fluctuations are minimal. In such systems, one can also identify MBL-like behavior (e.g., exponential-like dependence of transport properties on disordered strength $W$) [35, 36, 37, 38, 39, 40, 41, 42], being also realized experimentally in cold-atom systems [43, 44]. Recently, the debate of the MBL regime has been revived by the analytical study indicating that even at large disorder, the transport is not localized but rather subdiffusive [45], having the origin in insulating/blocking regions and related Griffiths effects already considered in MBL models [17, 46, 44].
@@ II Spin and energy diffusion
with $S=1/2$ spin operators and random fields $w_{l}\in[-W,W]$ with the uniform probability distribution. We consider the chain with $L$ sites and periodic boundary conditions numerically. To allow the Thouless approach to test transport properties via sensitivity to (twisted) boundary conditions, as well as the validity of RMT universality, [52, 34], we introduce in general a finite flux $\varphi\neq 0$. As with most MBL studies, we fix for convenience in numerical studies the (Ising) interaction corresponding to the isotropic case $\Delta=1$ and use $J=1$ as the unit of energy. Within the MBL problem, we are interested in the dynamical transport properties at finite temperature $T>0$, and as usual, for simplicity, we deal with the $T\to\infty$ limit. While d.c. transport quantities (i.e., the spin conductivity $\sigma^{0}_{s}$ and thermal conductivity $\kappa^{0}$) vanish at $T\to\infty$, the related diffusion constants do not and remain highly nontrivial quantities to determine.
@@
In the last equation, ${\cal D}_{s}(\omega)$ is expressed directly in terms of spin-current matrix elements among many-body (MB) eigenstates $|n\rangle$, $|m\rangle$, and $N_{st}$ representing the total number of MB states. The spin current $j_{s}$ and high-$T$ static susceptibility $\chi_{s}^{0}=\chi^{0}_{q\to 0}=\tilde{\chi}_{s}^{0}/T$ are within the considered model (1) given by
@@
where ${\cal D}_{s}^{0}={\cal D}_{s}(\omega\to 0)$ is d.c. spin diffusion constant and $\Delta e=\overline{e_{n+1}-e_{n}}$ is the average level spacing, or equivalently, $\rho=1/\Delta e$ the MB density of states. We note that the Thouless criterion, applied to the noninteracting Anderson model [52], requires $R<1$ for localization, but also that $R$ decreases/vanishes with increasing $L$. It is also evident that in the finite-$L$ system with $R<1$, one generally cannot expect the validity of RMT universality since the MB levels do not vary enough with $\varphi$ to satisfy the concept of avoided crossing. Consequently, the level sensitivity $R\sim 1$ is analogous to the criterion of merging the Heisenberg vs. Thouless time $\tau_{\mathrm{H}}\sim\tau_{\mathrm{T}h}$ [10], where $\tau_{\mathrm{H}}\sim 1/\Delta e$ and $\tau_{\mathrm{T}h}\sim L^{2}/{\cal D}^{0}_{s}$.
Figure 2: Level sensitivity $R$: (a) vs. $W$ for different $L$: for $L=14-18$ obtained via direct ED, and for $L=20-28$ via MCLM with approximated $\Delta e$ (see text for details), (b) the same averaged $R$ vs. $L$ for different $W$. The results for $L\leq 18$ ($L\geq 20$) were obtained as average over $N_{c}=10$ ($N_{c}=100$) random field configurations.
In Fig. 1, we present results for RMT markers for the spin current, $Y_{s}$, $Q_{s}$, and the energy current $Y_{e}$, $Q_{e}$, as well as the level sensitivity $R$, evaluated according to Eq. (9). Results are obtained via ED in the chain with $L=18$ sites for a few ($N_{c}=10$) fixed random configurations $\eta_{l}=[-1,1]$ with fields $w_{l}=W\eta_{l}$ steadily increasing with $W$. The sums over eigenstates in Eq. (3) are here performed over one-quarter of states, i.e., $\tilde{N}_{st}=N_{st}/4$ states in the middle of the MB spectrum. The presented results offer some nontrivial conclusions: (a) The variation of RMT markers vs. $W$ in Figs. 1(a,b) is quite similar for spin and energy current matrix elements. (b) The deviations for GOE values $Y_{s,e}\sim 1$ and $Q_{s,e}\sim 3$ appear in each sample quite simultaneously with the crossover at $R\sim 1$. (c) While the sample-sample fluctuations of d.c. values ${\cal D}^{0}_{s,e}$ are already substantial at intermediate $W\sim 1.5-2$ (see the discussion below), each sample still reveals $Q_{s,e}\sim 3$. This allows the conclusion that for fixed $w_{l}$ configurations, the variations of the calculated d.c. values within rather broad MB spectrum $e_{n}$ (one quarter of) are quite modest.
@@ IV Distribution of diffusion constants
It has been previously numerically established, at least for the spin diffusion ${\cal D}_{s}^{0}$ [32], that results reveal a broad distribution of values in reachable systems $L\lesssim 28$. The latter cannot be captured within a simple Gaussian distribution even at intermediate $W\sim 2$, becoming closer to the log-normal distribution up to the (here numerically) reachable $W^{*}\lesssim 4$. We note that such anomalous distributions are very distinct from the Anderson model of noninteracting particles in higher dimension $D>1$, where, due to large reachable sizes $L^{D}\sim 10^{7}$ [31] evidently yield already self-averaged results.
Figure 3: (a) Different presentations of the distribution of d.c. spin diffusion values ${\cal D}^{0}_{s}$: (1) vs. unrenormalized $W$, (2) using renormalized variance $\sum_{l}\eta^{2}_{l}/L=1/3$, and (3) presenting results vs. effective $\tilde{W}=W/(\bar{\eta}^{2}/3)$ (see text for details). (b) and (c) shows the distribution of results for ${\cal D}_{s}^{0}$ and ${\cal D}_{e}^{0}$, respectively, vs. effective $\tilde{W}$ obtained via MCLM for $L=24$, while the inset displays also the comparison with $L=20$ and $L=28$ results. Solid lines refer to the level sensitivity $R=1$, i.e., the estimated limit of validity of GOE. Dashed lines in panels (b,c) depict $\propto\exp(-bW/J)$ with $b=2.5$. The presented results are obtained from $N_{c}=100$ random-field samples.
@@
Figure 4: Cumulative distribution function (CDF) for (a) d.c. spin diffusion ${\cal D}_{s}^{0}$, (b) d.c. energy diffusion ${\cal D}_{e}^{0}$, obtained via MCLM on $L=24$ sites for different $W=0.5-4$, and (c) the direct comparison between both for $W=1.0,2.0,3.0,4.0$, including also $L=20\,,28$ results. The presented results are obtained from $N_{c}=100$ random-field samples.
@@ V Evidence for subdiffusion
Figure 5: System size $L$ dependence of spin diffusion ${\cal D}_{s}^{0}$ for $W=1,2,3$. Shown are MCLM (ED) results for $L\geq 20$ ($L\leq 18$) and $N_{c}=3$ different configurations. The dotted lines are guides to the eye, denoting the general trend with $L$.
@@
The above simple scenario for diffusion in MBL systems has implications. Namely, it should also apply to spin chains with quasiperiodic (QP) fields studied both theoretically/numerically [35, 36, 37, 38, 39, 40, 41, 42, 34], as well as experimentally [43, 44]. Indeed, similar exponential dependence on field strength $W$ has been found for ${\cal D}_{s}^{0}$ in QP systems [34]. With the essential difference that in the latter deterministic case (the absence of randomness), there are no significant (up to boundary effects) variations of ${\cal D}_{s}^{0}$ with actual field configuration. In contrast, for random but uncorrelated fields, as discussed in this study, there are large variations in $J_{l}^{\mathrm{eff}}$ resulting in variations of the effective distance between resonant islands $d$, as well as $w_{l}$ in the localized regions. These ingredients are the argument for the Griffiths scenario leading to domination and blocking of longer insulating islands [17, 46, luschen16, 32]. Furthermore, such arguments are also the basis for the analytical study supporting the subdiffusive transport in the thermodynamic limit [45].
Figure 6: Dynamical (a) spin and (b) energy diffusivity ${\cal D}_{s,e}(\omega)$ in the log-log scale, calculated for a single field configuration and $W=0.5-3.5$ on chains with different lengths, i.e., $L=18$ using ED, and $L=20-28$ employing MCLM for one random field sample.
@@
In contrast, one expects more pronounced effects in the dynamical diffusivity ${\cal D}_{s,e}(\omega)$ since results cover a much larger dynamical range, taking into account also $\omega>\delta\omega\sim 10^{-4}$. As the signature of subdiffusion one can identify the dependence ${\cal D}(\omega)\sim{\cal D}^{0}+c|\omega|^{\alpha}$ with $\alpha<1$ [17, 46] in the low $\omega<\omega^{*}$ range. Such behavior implies that even with vanishing (or very small) ${\cal D}^{0}$, the d.c. polarizability $\chi_{p}\propto\int\mathrm{d}\,\omega{\cal D}_{s}(\omega)/\omega^{2}$ [20] is diverging, in contrast to the localization defined by $\chi_{p}<\infty$ requiring $\alpha>1$. Numerical investigation of finite systems indicated that ${\cal D}_{s,e}(\omega)$ [and related conductivity $\sigma_{s}(\omega)$] at $\omega\sim{\cal O}(0.1)$ is indeed consistent with $\alpha\simeq 1$ [59, 16, 64, 19]. A similar conclusion (in a similar frequency range) can be reached by analysis of the short wave-length $q\to\pi$ response (i.e., the so-called imbalance) [25, 65, 32]. Here, contrary, we show that a closer analysis of $\omega\ll 1$ (obtained with upgraded MCLM method with $\delta\omega\sim 10^{-4}$ resolution) indicates that random-field results are generally at larger $W$ consistent with $\alpha<1$. The latter contrasts the QP field results, for which $\alpha\simeq 1$ is reported down to $\omega\to 0$.
@@
Figure 7: (a,b) Dynamical spin diffusivity ${\cal D}_{s}(\omega)-{\cal D}_{s}^{0}$ for two random field configurations with few $W$, calculated on $L=28$ chain. Each panel represents a different $w_{l}$ configuration. The black dashed lines correspond to the fits to the $c|\omega|^{\alpha}$ function in the $0.050<\omega<$ 0.05 region. Panel (c) depicts $\alpha$ dependence on the strength of the disorder $W$ (obtained from the fit).
@@
Figure 8: (a) Dynamical spin diffusivity ${\cal D}_{s}(\omega)$, calculated via MCLM for for quasiperiodic system ($L=27$). See text for details. (b) Frequency dependence of ${\cal D}_{s}(\omega)-{\cal D}_{s}^{0}$, together with fits to the $c|\omega|^{\alpha}$ function (black, dashed lines). The inset depicts $\alpha$ dependence on the (quasiperiodic) disorder strength.
In Fig. 6, we present numerical results for dynamical ${\cal D}_{s,e}(\omega)$, obtained for single field configurations but different $W=0.5-3.5$ on different system sizes $L=18-28$. While $L=18$ can be reached by ED directly evaluating Eq. (3), $L\geq 20$ results are obtained via MCLM with finite frequency resolution $\delta\omega\sim 10^{-4}$. Apart from fluctuations in MCLM results, being the consequence of finite $\delta\omega$ as well as the restricted sampling over initial microcanonical energies ${\cal E}$, results are consistent, provided that we input the same field configurations $w_{l}$. The dependence of MCLM results on specific $w_{l}$ configurations is presented in Fig. 7 for two different field configurations [with samples chosen from the middle of the distribution presented in Fig. 4(a)]. To focus on the frequency dependence we present ${\cal D}_{s}(\omega)-{\cal D}^{0}_{s}$, which allows for the analysis of $\propto|\omega|^{\alpha}$ fit in the $0.050<\omega<$ 0.05 region. It is evident from the results presented in Fig. 7(c), that $\alpha<1$ for $W\gtrsim 1$ and $\alpha\simeq 0.5$ for $W\simeq 2$ for both considered samples. Furthermore, our results indicate that although $\alpha<1$ for all considered $41<W<$ 4, its value slowly approaches $\alpha\sim 1$ for increasing $W$. Such linear frequency dependence of ${\cal D}_{s}(\omega)$ was previously reported [17, 64, 66, 25, 19, 30] in proximity to the (presumable) transition to MBL phase, $W>W_{c}$, based on the quantities averaged over random field configurations. Note that the true MBL phase reacquires ${\cal D}_{s}^{0}=0$ as well as $\alpha>1$. Furthermore, there are also some numerical evidence that characteristic subdiffusion scale $\omega^{*}$ is shrinking with disorder strength $W$, e.g., looking at Fig. 7a. The argument could be that $\omega^{*}$ should decrease together with strongly decreasing ${\cal D}_{s}^{0}$. Still, evidence is not conclusive, mostly due to numerical restrictions due to finite resolution $\delta\omega$ at large $W$.
Finally, we comment on the difference between the random-field and QP model. Note that the crossover to the MBL phase in the latter was found for $W\simeq 1.5$ [38, 34]. The QP field configuration is obtained with $w_{l}=W\cos(ql+\phi_{0})$, (nearly) irrational $q/(2\pi)=11/27\sim(3-\sqrt{5})/2$, and $\phi_{0}=0$. In Fig. 8(a), we show the full ${\cal D}_{s}(\omega)$ MCLM spectra (for $L=27$ sites), while in panel (b) we again focus on the frequency dependence of ${\cal D}_{s}(\omega)-{\cal D}^{0}_{s}$ with fits to $|\omega|^{\alpha}$. It is evident that the results for QP at low-$\omega$ are different: for all considered values $W$ we observe $\alpha\gtrsim 1$.